A pool owner is trying to fill his pool. The pump takes 6 hrs. A garden hose takes 10 hours. He wants to fill it up in a hurry. So he turns on the pump AND puts in the garden hose. Unfortunately, he leaves the drain open (which will drain the pool in 8 hours). How long will it take to fill the pool?
Answer by nerdybill(7384)

A pool owner is trying to fill his pool. The pump takes 6 hrs. A garden hose takes 10 hours. He wants to fill it up in a hurry. So he turns on the pump AND puts in the garden hose. Unfortunately, he leaves the drain open (which will drain the pool in 8 hours). How long will it take to fill the pool?
Let x = time (hours) it takes to fill pool
then
x(1/6 + 1/10 - 1/8) = 1
multiplying both sides by 6*10*8:
x(80 + 48 - 60) = 6*10*8
x(20 + 48) = 480
x(68) = 480
x = 480/68
x = 7.06 hours
or
x = 7 hours 3 minutes and 32 seconds